rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#207 — FS#4296 — DEC-iX

Attached to Project— Network
Maintenance
Whole Network
CLOSED
100%
All links towards the DEC-IX are down. More information are forthcoming.
Date:  Friday, 18 June 2010, 08:21AM
Reason for closing:  Done
Comment by OVH - Friday, 18 June 2010, 08:16AM

The corruption is due to an urgent maintenance of the switch on which we are linked to DEC-IX. The service has to be reset in ten minutes.


Comment by OVH - Friday, 18 June 2010, 08:18AM

The physical links are now reset but we are still isolated from lan peering at the moment.


Comment by OVH - Friday, 18 June 2010, 08:20AM

dec-ix faced some difficulties but bgp sessions are now forwarded.

Update 2010-06-18 05:35 UTC

We still have problems with sw021, all LACP port channels and some single ports did not
come up after the reboot and the switch forgot parts of the configuration. We restored
configuration and rebooted again, if this does not work we will revert to the old software.